OpenStudy (turtwig):

HELLPP! how do i "find the equation for y"? 5x - 7y = 10

OpenStudy (turtwig):

it's multiple choice

OpenStudy (igreen):

You have to get 'y' by itself on one side, start by subtracting 5x to both sides.

OpenStudy (turtwig):

-7y=10-5x ?

OpenStudy (igreen):

Yep! Now divide -7 to both sides.

OpenStudy (kiakat):

Yes, and then you can divide both sides by -7 to isolate y. Then you have the equation for y ;)

OpenStudy (turtwig):

mmm.... k so y=(how do i divide this side by -7?)

OpenStudy (turtwig):

@iGreen

OpenStudy (igreen):

\(-7y = 10 - 5x\) \(\dfrac{-7y}{-7} = \dfrac{10}{-7} = \dfrac{5x}{-7}\)

OpenStudy (igreen):

Sorry there should be a subtraction side on the right, not an equal sign.

OpenStudy (kiakat):

When you divide each side by -7, you're getting the y by itself while still keeping the -7 part of the equation

OpenStudy (igreen):

Yes, the -7's on the left cancel out, giving us: \(y = \dfrac{5}{7}x - \dfrac{10}{7}\)

OpenStudy (igreen):

\(\bbox [10pt, lime, border:5pt solid black]{\huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\) You can give medals by clicking 'Best Response'.

OpenStudy (turtwig):

<33